Find the remaining trigonometric functions of theta if: csc theta = 3 and theta terminates in Q II (#37)

37. csc B = 3 and e terminates in QI
Answer #1

we have that ,

\csc \theta=3

In second quadrant sin and cosec are positive , all others are negative .

we get other functions as

\sin \theta=\frac{1}{\csc \theta}=\frac{1}{3}

\cos \theta=\sqrt{1-\sin^2\theta}=\sqrt{1-\frac{1}{9}}=-\frac{\sqrt{8}}{3}

\sec \theta=\frac{1}{\cos \theta }=-\frac{3}{\sqrt{8}}

\tan \theta=\frac{\sin \theta}{\cos \theta }=\frac{1/3}{\sqrt{8}/3}=-\frac{1}{\sqrt{8}}

\cot \theta=\frac{1}{\tan \theta}=-\sqrt{8}